OpenStudy (anonymous):

lim cos(pi+h)+1 / h h->0 evaluate the limit

OpenStudy (anonymous):

lim tan(pi/4+h)-1/(h) someone please help me solve this h->0

OpenStudy (anonymous):

\[\cos(\lim_{h \rightarrow 0}(x+\pi))+\lim_{h \rightarrow 0}\frac{ 1 }{ h }=\cos(0+\pi)+\lim_{h \rightarrow 0}\frac{ 1 }{h }\] \[-1+\lim_{h \rightarrow0}\frac{ 1 }{ h}=-1+\infty=\]

OpenStudy (anonymous):

\[\infty\]

OpenStudy (anonymous):

@cheli320 in the second one:\[\tan(\frac{ \pi }{4+h })\] or\[\tan(\frac{ \pi }{ 4 }+h)\]

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

when you solve it do you get 2 as an answer?

OpenStudy (anonymous):

not yet.

OpenStudy (anonymous):

in the second question ,\[\frac{ \pi }{ 4 }+h\] or \[\frac{ \pi }{ 4+h }\]

OpenStudy (anonymous):

ok so when I do that and continue to work it out I get an answer of 2

OpenStudy (anonymous):

which form your question?

OpenStudy (anonymous):

first form

OpenStudy (anonymous):

\[\tan \frac{ \pi }{ 4 }-\frac{ 1 }{ 0}=-1+\infty=\infty\]

OpenStudy (anonymous):

thanks :)